# EquiReg: Equivariance Regularized Diffusion for Inverse Problems

### v1

@@ Appendix F Experiment Setup for PDE Reconstructions
Navier-Stokes equations. Following the methodology of [135], we model the time evolution of a vorticity field, $u(x,t)$, governed by:
@@
where ${\bm{w}}$ is the velocity field; $\nu=\frac{1}{1000}$, viscosity; and $f$, a fixed forcing term. The initial condition $a(x)$ is drawn from $\mathcal{N}(0,7^{3/2}(-\Delta+49\mathbf{I})^{-5/2})$ under periodic boundary conditions. The forcing term is $f(x)=0.1\left(\sin(2\pi(x_{1}+x_{2}))+\cos(2\pi(x_{1}+x_{2}))\right)$. We borrow the dataset from [51]. We note that this system has a reflection symmetry along the $x_{1}=x_{2}$ axis.
Implementation details. EquiReg, as a regularizer for diffusion posterior sampling, can be adapted to many inverse solvers in a plug-and-play manner. For PDE experiments, we use the same model weights and configurations as FunDPS [57]. Error rates are calculated using the $L^{2}$ relative error between the predicted and true solutions, averaged on 100 randomly selected test samples. We provide the information on the EquiReg scaling weights in Table 9.
@@ Appendix G Theoretical Analysis
In this section, we provide the proofs for Proposition 4.1 and Proposition 4.2.
G.1 Preliminary and Notations
@@
Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}\cup\{+\infty\}$ be a functional of probability distributions. The Wasserstein gradient flow of $\mathcal{F}$ is characterized by the minimizing movement scheme (also known as JKO scheme) introduced by [136]. For a fixed time step $\tau>0$, the sequence $(\rho_{k})_{k\in\mathbb{N}}$ of probability densities is defined recursively by:
@@
where $\bm{w}_{t}$ is standard Brownian motion, $\beta_{t}>0$ is a time-dependent variance schedule, and $f({\bm{x}},t)$ is a drift term. For instance, $f\equiv 0$ for a variance-exploding SDE and $f({\bm{x}},t)=-\frac{\beta_{t}}{2}{\bm{x}}$ for a variance-preserving SDE defined in [30]. In this work, we carry out our analysis under a more general setting.
@@
The corresponding evolution of probability density is
@@
$=-\nabla\cdot(\rho({\bm{x}},t)(-\nabla V({\bm{x}},t))+\alpha(t)\Delta\rho({\bm{x}},t),$ (26)
@@
Now we come back to Proposition 4.1. From eq. 14 we know that choosing
@@
in Lemma G.1 completes the proof, where $h$ is defined in Assumption G.1.
G.3 Detailed version of Proposition 4.2
@@
As discussed in the main text, since the conditional expectation is a linear combination over all possible values of ${\bm{x}}_{0}$, it may fall outside the data manifold, resulting in physically invalid samples. One of the central challenges in diffusion-based inverse sampling is guiding the sampling trajectory—generated by the reverse SDE dynamics—toward the data manifold. A common strategy is to incorporate regularization into the reverse SDE to encourage manifold adherence. In this work, building on the perspective of Wasserstein gradient flow as outlined above, we provide a novel interpretation of the role played by such regularization terms.
We show that the regularizer serves to reweight the contribution of different regions in the calculation of the underlying functional being minimized, $\Phi(\rho,t)$ defined in Proposition 4.1. Specifically, it amplifies the influence of regions where the density estimate is reliable (typically near the data manifold), while down-weighting regions with poor approximation quality of based on eq. 29, often corresponding to off-manifold samples.
[…]
@@
$\mathcal{R}({\bm{x}})\approx\begin{cases}0,&{\bm{x}}\in N(\mathcal{M})\\ \gg 1,&{\bm{x}}\text{ far away from }N(\mathcal{M})\end{cases}\quad\Rightarrow\quad\lambda({\bm{x}},t)\approx\begin{cases}1,&{\bm{x}}\in N(\mathcal{M})\\ 0,&{\bm{x}}\text{ far away from }N(\mathcal{M})\end{cases}.$
Next, we consider practical algorithms based on this reweighted functional. In practice, we only have the score function instead of the function value of $\log p_{T-t}({\bm{x}})$. Thus, the Wasserstein gradient flow associated with (33) is intractable since we cannot evaluate the weight function. We consider the following approximation based on $e^{\delta}\approx 1+\delta$,
@@
By Lemma G.1, the dynamics of ${\bm{x}}$ driven by the Wasserstein gradient flow associated with the approximated functional above is
@@ Appendix H Additional Background Information
Solving inverse problems with deep learning prior to diffusion models. Earlier works [137, 138, 139, 140] used deep neural networks as denoisers to solve inverse problems. Furthermore, deep generative models such as variational autoencoders (VAEs) [141], and generative adversarial networks (GANs) [142] were employed. Notable applications include compressed sensing [143] and MRI [144].
Resources for Definition 4.2 on vanishing-error autoencoders. Manifold constrained distribution-dependent equivariance error uses the notion of vanishing-error autoencoders [145, 146, 62] (Definition H.1), also known as an asymptotically-trained autoencoder [146] or a perfect autoencoder [62]. Vanishing-error autoencoders have previously been employed by diffusion-based inverse solvers to preserve the diffusion process on the manifold [62].
@@
A vanishing-error autoencoder under the manifold $\mathcal{M}$ with encoder $\mathcal{E}:\mathcal{X}\rightarrow\mathcal{Z}$ and decoder $\mathcal{D}:\mathcal{Z}\rightarrow\mathcal{X}$ with $\mathcal{Z}=\mathbb{R}^{k}$ where $k<d$, has zero reconstruction error under the support of the data distribution $\mathcal{X}$, i.e., $\forall{\bm{x}}\in\mathcal{X}\subset\mathcal{M}$, ${\bm{x}}=\mathcal{D}(\mathcal{E}({\bm{x}}))$. It follows that the decoder is surjective on the data manifold, $\mathcal{D}:\mathcal{Z}\rightarrow\mathcal{M}$ [62], and the encoder-decoder composition forms an identity map, i.e., $\forall{\bm{z}}\in\mathcal{M},{\bm{z}}=\mathcal{E}(\mathcal{D}({\bm{z}}))$.
@@
Equivariance. Let ${\bm{z}}\in\mathbb{R}^{d}$ and ${\bm{x}}=f({\bm{z}})\in\mathbb{R}^{d}$. For rotation and reflection equivariance, the transformations $T_{g}$ and $S_{g}$ can be defined by a rotation matrix $\bm{R}\in\mathbb{R}^{d\times d}$; then, a function $f$ with the rotation equivariant property would satisfy $\bm{R}{\bm{x}}=f(\bm{R}{\bm{z}})$. For translation equivariance, the transformations would be $T_{g}({\bm{z}})={\bm{z}}+g$ and $S_{g}({\bm{x}})={\bm{x}}+g$, where $g\in\mathbb{R}^{d}$. Hence, for a translation equivariance function $f$, we would have ${\bm{x}}+g=f({\bm{z}}+g)$. For the case where the output dimension is larger than the input, $f:\mathbb{R}^{k}\rightarrow\mathbb{R}^{d}$ with $d>k$, translation equivariance can be defined up to a discrete scale, i.e., $T_{g}({\bm{z}})={\bm{z}}+g$ and $S_{g}({\bm{x}})=T_{sg}({\bm{z}})$ where $s=\nicefrac{{d}}{{k}}$. The equivariance properties of translation, rotation, and reflections, combined, are referred to as E(3) symmetries. Without reflections, the symmetries form a Euclidean group SE(3) [112, 113].
Appendix I Computing Resources
@@
Appendix J Assets
Appendix K Broader Impacts
On the positive side, high-fidelity image restoration can improve downstream tasks in medical imaging, remote-sensing and environmental monitoring (e.g., denoising satellite observations to track pollution or deforestation). Likewise, accelerated PDE-solving via learned diffusion priors may enable faster, more accurate simulations for climate modeling, fluid-dynamics research, and engineering design. On the other hand, robust reconstruction methods could be misappropriated for privacy-invasive surveillance or to create deceptive imagery. We emphasize that our method does not amplify these existing risks.
@@
Our approach uses only publicly available datasets and standard pre-trained diffusion models, introducing no novel dual-use or privacy risks. Consequently, no additional safeguards are required.

### v2

@@ Appendix F Experiment Setup for PDE Reconstructions
Navier-Stokes equations. Following the methodology of (Li et al., 2020), we model the time evolution of a vorticity field, $u(x,t)$, governed by:
@@
where $\bm{w}$} is the velocity field; $\nu=\frac{1}{1000}$, viscosity; and $f$, a fixed forcing term. The initial condition $a(x)$ is drawn from $\mathcal{N}(0,7^{3/2}(-\Delta+49\mathbf{I})^{-5/2})$ under periodic boundary conditions. The forcing term is $f(x)=0.1\left(\sin(2\pi(x_{1}+x_{2}))+\cos(2\pi(x_{1}+x_{2}))\right)$. We borrow the dataset from (Huang et al., 2024). We note that this system has a reflection symmetry along the $x_{1}=x_{2}$ axis.
Implementation details. EquiReg, as a regularizer for diffusion posterior sampling, can be adapted to many inverse solvers in a plug-and-play manner. For PDE experiments, we use the same model weights and configurations as FunDPS (Yao et al., 2025). Error rates are calculated using the $L^{2}$ relative error between the predicted and true solutions, averaged on 100 randomly selected test samples. We provide the information on the EquiReg scaling weights in Table˜12.
@@ Appendix G Theoretical Analysis
G.1 Summary of the Analysis
The theoretical framework presented in this paper is intended to motivate and guide the design of effective regularizers. This perspective, grounded in optimal transport theory (Ferreira and Valencia-Guevara, 2018), serves as an intuitive interpretation of the dynamics and motivates the design of regularization strategies such as EquiReg. We note that whether diffusion models follow exact Wasserstein dynamics still remains an open problem (Zheng et al., 2025).
Proposition G.1.
Let $\rho({\bm{x}},t)$ be the distribution of ${\bm{x}}_{T-t}$ driven by the ideal reverse dynamics (eq.˜3). Then, the evolution of $\rho$ follows the Wasserstein-2 gradient flow associated with minimizing functional $\Phi(\rho,t)$ defined as $\beta_{T-t}\int[\rho\phi({\bm{x}},t)+\frac{1}{2}\rho\log\rho]d{\bm{x}}$, where $\phi({\bm{x}},t)=-(\log p_{T-t}({\bm{x}}|{\bm{y}})+\frac{1}{4}\|{\bm{x}}\|^{2})$.
Because the conditional expectation $\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{T-t}]$ is a linear combination of all candidate ${\bm{x}}_{0}$, the approximation remains relatively accurate when TtT-t is small (i.e., ${\bm{x}}_{T-t}$ stays close to the data manifold under low noise) but may incur high error for larger TtT-t, as shown in Figure˜1b. To mitigate this, we reweight the contributions to the first term of $\Phi$, down-weighting unreliable estimates, and amplifying the reliable ones. The resulting reweighted functional is
[…]
Proposition G.2.
(Informal) The evolution of $\rho$, the probability distribution of ${\bm{x}}_{T-t}$ driven by the practical and regularized reverse dynamics (eq.˜11), is an approximation of the Wasserstein-2 gradient flow associated with minimizing $\tilde{\Phi}$.
[…]
G.2 Preliminary and Notations
@@
Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}\cup\{+\infty\}$ be a functional of probability distributions. The Wasserstein gradient flow of $\mathcal{F}$ is characterized by the minimizing movement scheme (also known as JKO scheme) introduced by (Jordan et al., 1998). For a fixed time step $\tau>0$, the sequence $(\rho_{k})_{k\in\mathbb{N}}$ of probability densities is defined recursively by:
@@
where $\bm{w}_{t}$ is standard Brownian motion, $\beta_{t}>0$ is a time-dependent variance schedule, and $f({\bm{x}},t)$ is a drift term. For instance, $f\equiv 0$ for a variance-exploding SDE and $f({\bm{x}},t)=-\frac{\beta_{t}}{2}{\bm{x}}$ for a variance-preserving SDE defined in (Song et al., 2021). In this work, we carry out our analysis under a more general setting.
@@
We note that By Hörmander’s theorem, a smooth density $\rho({\bm{x}},t)$ exists for $t>0$, ensuring that the above $v$ is well-defined. The corresponding evolution of probability density is
@@
$=-\nabla\cdot(\rho({\bm{x}},t)(-\nabla V({\bm{x}},t))+\alpha(t)\Delta\rho({\bm{x}},t)),$ (27)
@@
Now we come back to Proposition˜G.1. From eq.˜15 we know that choosing
@@
in Lemma˜G.1 completes the proof, where $h$ is defined in Assumption˜G.1.
G.4 Detailed version of Proposition˜G.2
@@
As discussed in the main text, since the conditional expectation is a linear combination over all possible values of ${\bm{x}}_{0}$, it may fall outside the data manifold, resulting in physically invalid samples. One of the central challenges in diffusion-based inverse sampling is guiding the sampling trajectory, generated by the reverse SDE dynamics, toward the data manifold. A common strategy is to incorporate regularization into the reverse SDE to encourage manifold adherence. In this work, building on the perspective of Wasserstein gradient flow as outlined above, we provide a novel interpretation of the role played by such regularization terms.
We show that the regularizer serves to reweight the contribution of different regions in the calculation of the underlying functional being minimized, $\Phi(\rho,t)$ defined in Proposition˜G.1. Specifically, it amplifies the influence of regions where the density estimate is reliable (typically near the data manifold), while down-weighting regions with poor approximation quality of based on eq.˜30, often corresponding to off-manifold samples.
Following from what we have shown in the main text, $\Phi(\rho,t)$ has the form of $\beta_{T-t}\int[\rho\phi({\bm{x}},t)+\frac{1}{2}\rho\log\rho]d{\bm{x}}$ for a function $\phi({\bm{x}},t)$, which can be derived by (29). The $\log p_{t}({\bm{y}}|{\bm{x}})$ term in (29) or $\nabla\log p_{t}({\bm{y}}|{\bm{x}})$ term in (28), equivalently, is computed based on approximation (30). We denote the corresponding approximation of $\phi({\bm{x}},t)$ as $\hat{\phi}({\bm{x}},t)$. As discussed in the main text, we can assume without loss of generality that $\phi({\bm{x}},t)<0$ and $\hat{\phi}({\bm{x}},t)<0$. We have
@@
[…]
Next, we consider practical algorithms based on this reweighted functional. In practice, we only have the score function instead of the function value of $\log p_{T-t}({\bm{x}})$. Thus, the Wasserstein gradient flow associated with (34) is intractable since we cannot evaluate the weight function. We consider the following approximation based on $e^{\delta}\approx 1+\delta$ when $\delta$ is sufficiently small,
@@
By Lemma˜G.1, the dynamics of $\bm{x}$} driven by the Wasserstein gradient flow associated with the approximated functional above is
@@ Appendix H Additional Background Information
Solving inverse problems with deep learning prior to diffusion models. Earlier works (Metzler et al., 2016; Romano et al., 2017; Zhang et al., 2017; Metzler et al., 2017) used deep neural networks as denoisers to solve inverse problems. Furthermore, deep generative models such as variational autoencoders (VAEs) (Kingma, 2013), and generative adversarial networks (GANs) (Goodfellow et al., 2014) were employed. Notable applications include compressed sensing (Bora et al., 2017) and MRI (Jalal et al., 2021).
Applications on diffusion models to solve inverse problems. Most popular applications include image restoration (Chung et al., 2023; 2022b; Kawar et al., 2022; Lugmayr et al., 2022; Saharia et al., 2022; Song et al., 2023a; Rout et al., 2023; Zhu et al., 2023; Zhang et al., 2025a; Zirvi et al., 2025), medical imaging (Song et al., 2022; Chung and Ye, 2022; Chung et al., 2022a; Hung et al., 2023; Dorjsembe et al., 2024; Li et al., 2024; Kazerouni et al., 2023; Bian et al., 2024), and solving partial differential equations (PDEs) (Isakov, 2006; Huang et al., 2024; Shysheya et al., 2024; Liu et al., 2023; Li et al., 2025; Baldassari et al., 2023; Mammadov et al., 2024a; Yao et al., 2025). On the methodology side, there has been numerous advancements (Chung et al., 2023; 2022b; Kawar et al., 2022; Lugmayr et al., 2022; Saharia et al., 2022; Song et al., 2023a; Rout et al., 2023; Zhu et al., 2023; Zhang et al., 2025a; Zirvi et al., 2025; Song et al., 2022; Chung and Ye, 2022; Chung et al., 2022a; Hung et al., 2023; Dorjsembe et al., 2024; Li et al., 2024; Kazerouni et al., 2023; Bian et al., 2024; Huang et al., 2024; Shysheya et al., 2024; Mammadov et al., 2024b; Cardoso et al., 2024).
Resources for Definition˜3.2 on vanishing-error autoencoders. Manifold constrained distribution-dependent equivariance error uses the notion of vanishing-error autoencoders (Shao et al., 2018; Anders et al., 2020; He et al., 2024) (Definition˜H.1), also known as an asymptotically-trained autoencoder (Anders et al., 2020) or a perfect autoencoder (He et al., 2024). Vanishing-error autoencoders have previously been employed by diffusion-based inverse solvers to preserve the diffusion process on the manifold (He et al., 2024).
@@
A vanishing-error autoencoder under the manifold $\mathcal{M}$ with encoder $\mathcal{E}:\mathcal{X}\rightarrow\mathcal{Z}$ and decoder $\mathcal{D}:\mathcal{Z}\rightarrow\mathcal{X}$ with $\mathcal{Z}=\mathbb{R}^{k}$ where $k<d$, has zero reconstruction error under the support of the data distribution $\mathcal{X}$, i.e., $\forall{\bm{x}}\in\mathcal{X}\subset\mathcal{M}$, ${\bm{x}}=\mathcal{D}(\mathcal{E}({\bm{x}}))$. It follows that the decoder is surjective on the data manifold, $\mathcal{D}:\mathcal{Z}\rightarrow\mathcal{M}$ (He et al., 2024), and the encoder-decoder composition forms an identity map, i.e., $\forall{\bm{z}}\in\mathcal{M},{\bm{z}}=\mathcal{E}(\mathcal{D}({\bm{z}}))$.
@@
Equivariance. Let ${\bm{z}}\in\mathbb{R}^{d}$ and ${\bm{x}}=f({\bm{z}})\in\mathbb{R}^{d}$. For rotation and reflection equivariance, the transformations $T_{g}$ and $S_{g}$ can be defined by a rotation matrix $\bm{R}\in\mathbb{R}^{d\times d}$; then, a function $f$ with the rotation equivariant property would satisfy $\bm{R}{\bm{x}}=f(\bm{R}{\bm{z}})$. For translation equivariance, the transformations would be $T_{g}({\bm{z}})={\bm{z}}+g$ and $S_{g}({\bm{x}})={\bm{x}}+g$, where $g\in\mathbb{R}^{d}$. Hence, for a translation equivariance function $f$, we would have ${\bm{x}}+g=f({\bm{z}}+g)$. For the case where the output dimension is larger than the input, $f:\mathbb{R}^{k}\rightarrow\mathbb{R}^{d}$ with $d>k$, translation equivariance can be defined up to a discrete scale, i.e., $T_{g}({\bm{z}})={\bm{z}}+g$ and $S_{g}({\bm{x}})=T_{sg}({\bm{z}})$ where $s=\nicefrac{{d}}{{k}}$. The equivariance properties of translation, rotation, and reflections, combined, are referred to as E(3) symmetries. Without reflections, the symmetries form a Euclidean group SE(3) (Thomas et al., 2018; Fuchs et al., 2020).
E(3), SE(3), and SO(3) are important symmetry groups in 3D Euclidean space, with well-established applications in physics and chemistry, computer vision, and reinforcement learning (Cohen and Welling, 2016; Thomas et al., 2018; Hoogeboom et al., 2022; Xu et al., 2024; Park et al., 2025). Finally, our contributions are complementary to, and can be combined with, the growing literature on meta-learning and automatic symmetry discovery to learn symmetry groups and their actions directly from data (Zhou et al., 2021; Quessard et al., 2020; Dehmamy et al., 2021; Mohapatra et al., 2025).
Data manifold hypothesis. Let data ${\bm{x}}\in\mathcal{X}\subset\mathbb{R}^{d}$ be in an ambient space of dimension $d$ with support $\mathcal{X}$ distribution. We assume that data are sampled from a low-dimensional manifold $\mathcal{M}$ (Cayton and others, 2005; Ma and Fu, 2012) embedded in a high-dimensional space (Assumption˜H.1). This hypothesis is popular in machine learning (Bordt et al., 2023), and has been studied mathematically in the literature (Narayanan and Mitter, 2010; Bortoli, 2022). Moreover, empirical evidence in image processing supports the manifold hypothesis (Weinberger and Saul, 2006; Fefferman et al., 2016), and diffusion-based solvers assume this property (He et al., 2024; Chung et al., 2022b; 2023).
Assumption H.1 (Manifold Hypothesis).
Appendix I Additional figures on MPE functions
Table 13: DPS superresolution with $\lambda=0.01$ using different MPE functions.
(a) FFHQ 256.
MPE function PSNR SSIM LPIPS FID
None 23.160 (1.923) 0.657 (0.072) 0.193 (0.057) 129.528
LDM Encoder (FFHQ) 26.581 (2.457) 0.773 (0.044) 0.120 (0.030) 87.437
CNN Autoencoder (FFHQ) 26.866 (1.943) 0.771 (0.044) 0.116 (0.029) 85.352
Pretrained ResNet50 26.873 (1.941) 0.771 (0.044) 0.116 (0.029) 85.138
Pretrained CLIP 26.860 (1.942) 0.771 (0.044) 0.116 (0.029) 85.495
(b) ImageNet.
MPE function PSNR SSIM LPIPS FID
None 19.727 (4.292) 0.407 (0.180) 0.541 (0.182) 446.829
LDM Encoder (ImageNet) 22.200 (4.295) 0.568 (0.146) 0.384 (0.130) 311.636
CNN Autoencoder (ImageNet) 22.178 (4.294) 0.568 (0.148) 0.375 (0.125) 312.530
Pretrained ResNet50 22.176 (4.290) 0.568 (0.148) 0.375 (0.125) 314.590
Pretrained CLIP 22.177 (4.293) 0.568 (0.148) 0.376 (0.125) 313.468
(a) FFHQ 256.
(b) ImageNet.
Figure 19: Equivariance error vs. $\sigma$ noise added. As more noise is added, equivariance error, computed with all MPE functions, increases.
Appendix J Computing Resources
@@ Appendix I Additional figures on MPE functions
Appendix K Assets
@@
Our approach uses only publicly available datasets and standard pre‑trained diffusion models, introducing no novel dual‑use or privacy risks. Consequently, no additional safeguards are required.